OpenStudy (rainbow_dash):

Solve 27^2/3. I know how to this but i completely forgot and a simple explanation would help.

OpenStudy (anonymous):

27 = 3^3 x^(a/b) = (x^(1/b))^a 27^(2/3)=(27^(1/3))^2 = ((3^3)^(1/3))^2=3^2=9

OpenStudy (rainbow_dash):

oh dang.

OpenStudy (rainbow_dash):

thank you